Show all steps in obtaining the analytical solution of dU 6 - 4e-3 with U(0)=0 dt
Answer #1

e-3t te Colution du dt -=4e-3t with Ulo)=0 du dt e 4/6 - -3+ integrating both side wortit e-3t de s du = lys ) ч/ С fo -3 បុ